Sean is driving to visit his friend in another state. In the first hour, he travels 1/5 of the total distance. In the second hour, he travels 34 miles. If he has 194 more miles to go, how far apart do the friends live?
Answer by mananth(16946)

let distance be x
1/5 *x+34+194=x
4x/5= 228
x=5*228/4
x=5*57
x=285 miles